OpenStudy (anonymous):

write in radical form. a. (3x)^1/2 b. 3x^1/2 c. 7^-1/2

OpenStudy (saifoo.khan):

\[\sqrt{3x}\] \[3 \sqrt x\] \[\frac{1}{\sqrt 7}\]

OpenStudy (anonymous):

do you mind explaining to me how you got those answer? pleaase

OpenStudy (saifoo.khan):

anything raise to the power 1/2 means square root. so, (3x)^1/2 means the whole thing square root. 3 just x ^1/2 means 3 and then square root x. while in the last one, it has negative expoenent. so we will make it positive by flipping, 1/7^(1/2) now same as before, 1/ sqrt(7)

OpenStudy (anonymous):

OOOOH okayy. that makes sense ! thank you ! (:

OpenStudy (saifoo.khan):

you welcome. please hit good answer.

OpenStudy (saifoo.khan):

ty. [=

OpenStudy (anonymous):

now, i need help on another question xD Write using exponents. a. \[\sqrt[5]{a ^{2}}\] b. \[(\sqrt[3]{b})^{2}\] c. \[\sqrt[4]{5t ^{3}}\] d. \[1/\sqrt[5]{x ^{4}}\]